# Kernel Ridge Riesz Representers: Generalization Error and Mis-specification

### v3

@@ Abstract
Kernel balancing weights provide confidence intervals for average treatment effects, based on the idea of balancing covariates for the treated group and untreated group in feature space, often with ridge regularization. Previous works on the classical kernel ridge balancing weights have certain limitations: (i) not articulating generalization error for the balancing weights, (ii) typically requiring correct specification of features, and (iii) providing inference for only average effects.
I interpret kernel balancing weights as kernel ridge Riesz representers (KRRR) and address these limitations via a new characterization of the counterfactual effective dimension. KRRR is an exact generalization of kernel ridge regression and kernel ridge balancing weights. I prove strong properties similar to kernel ridge regression: population $L_{2}$ rates controlling generalization error, and a standalone closed form solution that can interpolate. The framework relaxes the stringent assumption that the underlying regression model is correctly specified by the features. It extends inference beyond average effects to heterogeneous effects, i.e. causal functions. I use KRRR to infer heterogeneous treatment effects, by age, of 401(k) eligibility on assets.
@@ 1 Introduction
Kernel balancing weights adapt kernel methods to conduct inference on average effects, e.g. the average treatment effect when the treatment is randomly assigned conditional on covariates. [^1] Kernel ridge balancing weights for inference also directly apply to the average treatment effects for the treated and untreated subpopulations, as well as subpopulations defined in terms of a discrete covariate taking a finite number of values. The central idea is to balance the feature-mapped covariates in the treated group with those in the untreated group, subject to regularization. The weights that achieve such balance, when applied to the observed outcomes, estimate the average treatment effect and appear in its confidence interval. These weights are widely used in social science via popular R packages, and they are classical in statistics [Speckman, 1979].
The motivation for kernel balancing weights is to inherit the simplicity and flexibility of kernel methods. However, not all of the familiar properties of, say, kernel ridge regression have been fully articulated for classical kernel ridge balancing weights, e.g. population $L_{2}$ rates of convergence for generalization error, a simple standalone closed form solution, and robustness to mis-specification of features. Such properties would lead to strong semiparametric guarantees for causal parameters via well-known “targeting” and “debiasing” arguments, which tolerate some types of mis-specification and provide inference for causal functions.
This paper’s main contribution is a population $L_{2}$ rate controlling the generalization error of the classical kernel ridge balancing weights previously formulated in the literature. The rate coincides with the lower bound of regression in the reproducing kernel Hilbert space (RKHS) framework. It holds under mis-specification of features. A technical innovation powers the result: I define the counterfactual effective dimension, and prove it is not much greater than the actual effective dimension. This technique may be of independent interest.
@@
Conceptually, I study the kernel ridge Riesz representer (KRRR) as an exact generalization of kernel ridge regression and kernel ridge balancing weights. This generalization formalizes the sense in which the population $L_{2}$ rate is sometimes optimal. For treatment effects, I verify inference with rate-optimal regularization, rather than undersmoothing, of the ridge penalty. The generalization also allows kernel ridge balancing weights to conduct inference on causal functions, e.g. heterogeneous policy effects and heterogeneous treatment effects. [^2] Heterogeneous treatment effects are a function of a continuous covariate; see Section 7.
Section 2 situates these contributions within the context of related work. Section 3 formalizes the key assumptions and summarizes the main results. Section 4 describes the procedure, derives its standalone closed form, and clarifies how it extends known procedures. Section 5 theoretically justifies the procedure with population $L_{2}$ rates of generalization error, which imply strong semiparametric guarantees. Section 6 proves the rates via the counterfactual effective dimension. Section 7 demonstrates nominal coverage in heterogeneous treatment effect simulations, with less bias than some previous estimators, then infers the heterogeneous treatment effects of 401(k) eligibility on assets as a function of age. The effect appears to be five times stronger for 45-60 year olds, compared to 30 year olds. Section 8 concludes.
@@ 2 Related work
I extend the conceptual framework of kernel balancing weights [Speckman, 1979, Hazlett, 2020, Wong and Chan, 2018, Zhao, 2019, Kallus, 2020, Hirshberg and Wager, 2021, Hirshberg et al., 2019, Chernozhukov et al., 2020]. Works in this literature often focus on kernel ridge balancing weights as a vector that possibly converges in $\ell_{2}$ [Hirshberg and Wager, 2021, Hirshberg et al., 2019], whereas I will study them as evaluations of a function that converges in population $L_{2}$. The former is the training error for a balancing weight that is defined in-sample; see Corollary 4.3. By contrast, I analyze generalization error for a balancing weight that can interpolate and that can be evaluated out-of-sample. This distinction is important for stronger semiparametric conclusions that allow more general causal parameters, sample splitting, and regression models beyond the RKHS; see Remark 5.3.
@@
[Bruns-Smith et al., 2023] prove equivalences among various formulations of the classical kernel ridge balancing weights, including a minimax formulation with ridge regularization. Regardless of nomenclature, many of the salient properties of these classical balancing weights, such as population $L_{2}$ rates of generalization error and a standalone closed form, appear to be contributions. I recover the properties that were known, e.g. the nonasymptotic balance-variance trade-off.
@@
The properties that I prove for KRRR generalize known results for kernel ridge regression. I obtain $L_{2}$ rates using integral operator techniques for kernel methods, drawing on [Smale and Zhou, 2007, Caponnetto and De Vito, 2007, Fischer and Steinwart, 2020], among others. See [Fischer and Steinwart, 2020] for references and a review of integral operator techniques and empirical process techniques used to analyze RKHS estimators in various norms, many of which rely on the effective dimension. My characterization of the counterfactual effective dimension appears to be new, and it avoids auxiliary approximation assumptions. The closed form solution of KRRR echoes the classical solutions of [Kimeldorf and Wahba, 1971, Schölkopf et al., 2001], and it builds on the techniques of [Chernozhukov et al., 2020].
@@
I also build on the insight from semiparametric theory that the balancing weights for the average treatment effect are a special case of Riesz representers for causal parameters, which can be directly estimated [Robins et al., 2007, Avagyan and Vansteelandt, 2021, Chernozhukov et al., 2022d, Hirshberg and Wager, 2021, Chernozhukov et al., 2022c, Smucler et al., 2019, Hirshberg et al., 2019, Chernozhukov et al., 2020]. Similar to [Chernozhukov et al., 2022d, Chernozhukov et al., 2022c, Chernozhukov et al., 2020], I provide population $L_{2}$ rates. A key point of departure is that I study KRRR, which is a different estimator more closely aligned with kernel ridge regression, classical kernel ridge balancing weights, and popular R packages. Another distinction is that I place standard RKHS assumptions, which facilitates comparison with known optimality theory, including rate-optimal regularization. Finallly, I avoid some auxiliary approximation assumptions of those works by analyzing the counterfactual effective dimension; see Remark 6.1.
@@
For inference on the causal parameter, the contribution of this work is not to provide new semiparametric theory. Rather, it is to demonstrate that KRRR has good properties similar to kernel ridge regression that were not fully articulated, and hence it is compatible with strong existing results, e.g. [Zheng and van der Laan, 2011, Kennedy et al., 2017, Chernozhukov et al., 2018, van der Laan et al., 2018, Chernozhukov et al., 2022a, Rotnitzky et al., 2021, Kennedy, 2023, Chernozhukov et al., 2023], among many others. These strong results allow for some mis-specification and rate-optimal regularization, rather than the correct specification and undersmoothing studied by e.g. [Hirshberg et al., 2019, Mou et al., 2023]. See e.g. [Hirshberg and Wager, 2021, Ben-Michael et al., 2021] for detailed citations on balancing weights outside of the RKHS. See e.g. [van der Laan and Robins, 2003, van der Laan and Rose, 2011, Chernozhukov et al., 2022a] for detailed citations on classical and modern semiparametric theory.
@@
This paper’s initial draft was circulated in February 2021 [Singh, 2021]. Prior to the current draft, [Wang et al., 2021] extend the kernel balancing weights of [Wong and Chan, 2018] to heterogeneous treatment effects and prove consistency under correct specification, without formal inference, a closed form, or $L_{2}$ rates for balancing weights. This paper’s results are complementary: formal inference, some mis-specification, a closed form, and $L_{2}$ rates for balancing weights, within a framework that includes other estimands. Subsequent works extend kernel balancing weights to confounding bridges [Kallus et al., 2021, Ghassami et al., 2021] and study other nonlinear function spaces, e.g. neural networks [Chernozhukov et al., 2021, Chernozhukov et al., 2022b].
@@ 3.1 Warm up: Balancing weights
These two formulations suggest estimation approaches. One option is to estimate the regression as $\hat{\gamma}$ then the ATE as $\mathbb{E}_{n}\{\hat{\gamma}(1,X)-\hat{\gamma}(0,X)\}$, where $\mathbb{E}_{n}(\cdot)$ is the empirical mean. Another is to estimate the balancing weight as $\hat{\alpha}$ then the ATE as $\mathbb{E}_{n}\{Y\hat{\alpha}(D,X)\}$. One may also combine these estimators into a so-called doubly robust procedure $\mathbb{E}_{n}[\hat{\gamma}(1,X)-\hat{\gamma}(0,X)+\hat{\alpha}(D,X)\{Y-\hat{\gamma}(D,X)\}]$ [Robins et al., 1995, Robins and Rotnitzky, 1995]. Each approach also has a sample splitting variant [Bickel, 1982, Schick, 1986, Klaassen, 1987], where $(\hat{\gamma},\hat{\alpha})$ are estimated from, say $n/2$ held out observations, and the empirical mean is taken with respect to the remaining $n/2$ observations. For the sample splitting variant, $\hat{\alpha}$ must be a function that can be evaluated out-of-sample.
@@
Previous work proposes a kernel ridge balancing weight $\tilde{\alpha}$. My main contribution is to analyze its generalization error $\|\tilde{\alpha}-\alpha_{0}\|_{2}$, achieving the well-known lower bound for regression in the RKHS framework, without auxiliary approximation assumptions. Once a population $L_{2}$ rate is articulated, general semiparametric guarantees from the literature become available. These guarantees allow inference for causal functions and some mis-specification, which previous works on the classical kernel ridge balancing weights typically disallow.
@@ 3.2 Generalizing from balancing weights to Riesz representers
The ATE is a special case of a causal parameter. This paper considers parameters of the form $\theta_{0}=\mathbb{E}\{m(W,\gamma_{0})\}$, where $W$ concatenates observed variables, $\gamma_{0}(W)=\mathbb{E}(Y|W)$ is a regression, and the formula $m$ accommodates various parameters beyond the ATE. For ATE, $W=(D,X)$ and $m(w,f)=f(1,x)-f(0,x)$, but $W$ and $m$ may be more general. I study a class defined by a regularity condition generalizing the idea that the propensity score is bounded away from zero and one.
Assumption 3.1 (Mean square continuity).
@@
Nonasymptotic analysis allows $\bar{M}$ to be a diverging sequence as the sample size increases. In particular, this class includes functionals of the form $f\mapsto\mathbb{E}\{m_{h}(W,f)\}$ where $m_{h}(w,f)=\ell_{h}(w)\check{m}(w,f)$. The weighting $\ell_{h}(w)$ is indexed by a vanishing bandwidth $h\downarrow 0$, which also indexes $\bar{M}=\bar{M}_{h}\uparrow\infty$. In this thought experiment, a sequence of semiparametric quantities, each satisfying Assumption 3.1, provides a pointwise approximation to a nonparametric causal function evaluated at a point. To be concrete, I provide two examples.
Example 3.1 (Heterogeneous policy effects).
Here $W=X$ are regressors and $\check{m}(w,f)=f\{t(x)\}-f(x)$, where $t$ is a known mapping that serves as the counterfactual policy of interest [Stock, 1989]. The weighting $\ell_{h}(w)=\textsc{local}\{(x_{1}-x_{1}^{*})/h\}$ is a localization around the location $X_{1}=x_{1}^{*}$ (see Appendix A for details). Suppose that the density ratio is bounded above, i.e. $\omega(W)=\textsc{density}\{t(W)\}/\textsc{density}(W)\leq\bar{\omega}$ almost surely. Suppose that the weighting function at a fixed $h$ is bounded, i.e. $|\ell_{h}(W)|\leq\bar{\ell}_{h}$ almost surely. Then $\bar{M}_{h}=2\bar{\ell}_{h}^{2}(\bar{\omega}+1)$.
Example 3.2 (Heterogeneous treatment effects).
@@
For further interpretation, suppose the the first covariate $X_{1}$ is continuous. In the finite sample, Example 3.2 is the heterogeneous treatment effect for the subpopulation whose first covariate is local to $x_{1}^{*}$. As $h\downarrow 0$, Example 3.2 converges to the heterogeneous treatment effect for the subpopulation whose first covariate value is $x_{1}^{*}$. Section 7 gives a real world example where formal inference leads to economic insights.
Taking $\ell_{h}(w)=1$, Examples 3.1 and 3.2 reduce to the average policy effect and average treatment effect, respectively. The class defined by Assumption 3.1 contains not only average effects but also heterogeneous effects, i.e. pointwise approximations of causal functions.
The balancing weight is a special case of a Riesz representer. It is well known that each parameter within the class defined by Assumption 3.1 has both a regression formulation and a generalized balancing weight formulation.
Lemma 3.1 (Riesz representation; Lemma S3.1 of [Chernozhukov et al., 2022d]).
Suppose Assumption 3.1 holds and $\gamma_{0}\in\mathcal{G}\subset L_{2}$. Then there exists a Riesz representer $\alpha_{0}\in L_{2}$ such that $\mathbb{E}\{m(W,f)\}=\mathbb{E}\{\alpha_{0}(W)f(W)\}$ for all $f\in\mathcal{G}$. Moreover, there exists a unique minimal Riesz representer $\alpha_{0}^{\min}\in closure\{span(\mathcal{G})\}$ that satisfies this equation.
For the rest of the paper, I will set aside the “correct specification” assumption that $\gamma_{0}\in\mathcal{G}$, where $\mathcal{G}$ is a known subset of $L_{2}$ that can be imposed in estimation of $\hat{\gamma}$. As such, there exists a unique Riesz representer $\alpha_{0}=\alpha_{0}^{\min}$. See Remark 5.3 for a list of many works that place “correct specification” style assumptions requiring $\gamma_{0}$ to be within the RKHS.
For a functional of the form $f\mapsto\mathbb{E}\{m_{h}(W,f)\}$ where $m_{h}(w,f)=\ell_{h}(w)\check{m}(w,f)$, its Riesz representer takes the form $\alpha_{h}(w)=\ell_{h}(w)\check{\alpha}_{0}(w)$, where $\check{\alpha}_{0}(w)$ is the Riesz representer for $f\mapsto\mathbb{E}\{\check{m}(W,f)\}$.
In summary, previous works on inference with classical kernel ridge balancing weights focus on a limited class of causal parameters, namely average effects. They also typically require correct specification of $\gamma_{0}$ within the RKHS. I study the broader class defined in Assumption 3.1, which includes heterogeneous policy effects (Example 3.1) and heterogeneous treatment effects (Example 3.2). I take the more agnostic stance that $\gamma_{0}\in L_{2}$. By analyznig generalization error $\|\tilde{\alpha}-\alpha_{0}\|_{2}$, general semiparametric guarantees become immediate.

### v4

# Kernel Ridge Riesz Representers: Generalization, Mis-specification, and the Counterfactual Effective Dimension

@@ Abstract
Kernel balancing weights provide confidence intervals for average treatment effects, based on the idea of balancing covariates for the treated group and untreated group in feature space, often with ridge regularization. Previous works on the classical kernel ridge balancing weights have certain limitations: (i) not articulating generalization error for the balancing weights, (ii) typically requiring correct specification of features, and (iii) justifying Gaussian approximation for only average effects.
I interpret kernel balancing weights as kernel ridge Riesz representers (KRRR) and address these limitations via a new characterization of the counterfactual effective dimension. KRRR is an exact generalization of kernel ridge regression and kernel ridge balancing weights. I prove strong properties similar to kernel ridge regression: population $L_{2}$ rates controlling generalization error, and a standalone closed form solution that can interpolate. The framework relaxes the stringent assumption that the underlying regression model is correctly specified by the features. It extends Gaussian approximation beyond average effects to heterogeneous effects, justifying confidence sets for causal functions. I use KRRR to quantify uncertainty for heterogeneous treatment effects, by age, of 401(k) eligibility on assets.
@@ 1   Introduction
Kernel balancing weights adapt kernel methods to quantify uncertainty on average effects, e.g. the average treatment effect when the treatment is randomly assigned conditional on covariates. [^1] Kernel ridge balancing weights also apply to confidence intervals for the average treatment effects on the treated and untreated subpopulations, as well as subpopulations defined in terms of a discrete covariate taking a finite number of values. This paper justifies confidence sets for causal functions. The central idea is to balance the feature-mapped covariates in the treated group with those in the untreated group, subject to regularization. The weights that achieve such balance, when applied to the observed outcomes, estimate the average treatment effect and appear in its confidence interval. These weights are widely used in social science via popular R packages, and they are classical in statistics [Speckman, 1979].
The motivation for kernel balancing weights is to inherit the simplicity and flexibility of kernel methods. However, not all of the familiar properties of, say, kernel ridge regression have been fully articulated for classical kernel ridge balancing weights, e.g. population $L_{2}$ rates of convergence for generalization error, a simple standalone closed form solution, and robustness to mis-specification of features. Such properties would lead to strong semiparametric guarantees for causal parameters via well-known “targeting” and “debiasing” arguments, which tolerate some types of mis-specification and which provide Gaussian approximation for causal functions, justifying confidence sets with nominal coverage.
This paper’s main contribution is a population $L_{2}$ rate controlling the generalization error of the classical kernel ridge balancing weights. The rate coincides with the lower bound of regression in the reproducing kernel Hilbert space (RKHS) framework. It holds under mis-specification of features. A technical innovation powers the result: I define the counterfactual effective dimension, and prove it is not much greater than the actual effective dimension. This technique may be of independent interest.
@@
Conceptually, I study the kernel ridge Riesz representer (KRRR) as an exact generalization of kernel ridge regression and kernel ridge balancing weights. This generalization formalizes the sense in which the population $L_{2}$ rate is sometimes optimal. For treatment effects, I verify inference with rate-optimal regularization, rather than undersmoothing, of the ridge penalty. The generalization also justifies the construction of confidence sets for causal functions, e.g. heterogeneous policy effects and heterogeneous treatment effects, from kernel ridge balancing weights. [^2] Heterogeneous treatment effects are a function of a continuous covariate; see Section 7.
Section 2 situates these contributions within the context of related work. Section 3 formalizes the key assumptions. Section 4 describes the procedure, derives its standalone closed form, and clarifies how it extends known procedures. Section 5 theoretically justifies the procedure with population $L_{2}$ rates of generalization error, which imply strong semiparametric guarantees. Section 6 proves the rates via the counterfactual effective dimension. Section 7 demonstrates nominal coverage in heterogeneous treatment effect simulations, with less bias than some previous estimators, then quantifies uncertainty for the heterogeneous treatment effects of 401(k) eligibility on assets as a function of age. The effect appears to be five times stronger for 45-60 year olds, compared to 30 year olds.
@@ 2   Related work
I extend the conceptual framework of kernel balancing weights [Speckman, 1979, Hazlett, 2020, Wong and Chan, 2018, Zhao, 2019, Kallus, 2020, Hirshberg and Wager, 2021, Hirshberg et al., 2019, Chernozhukov et al., 2020]. Works in this literature often focus on kernel ridge balancing weights as a vector that possibly converges in $\ell_{2}$ [Hirshberg and Wager, 2021, Hirshberg et al., 2019], whereas I will study them as evaluations of a function that converges in population $L_{2}$. The former is the training error for a balancing weight that is defined in-sample; see Corollary 4. By contrast, I analyze generalization error for a balancing weight that can interpolate and that can be evaluated out-of-sample. This distinction is important for stronger semiparametric conclusions: coverage guarantees for causal functions, sample splitting, and regression models beyond the RKHS; see Remark 3.
@@
[Bruns-Smith et al., 2023] prove equivalences among various formulations of the classical kernel ridge balancing weights, including a minimax formulation with ridge regularization. Regardless of nomenclature, many of the salient properties of these classical balancing weights, such as population $L_{2}$ rates of generalization error and a standalone closed form, appear to be contributions. I recover the known, nonasymptotic balance-variance trade-off.
@@
The properties that I prove for KRRR generalize known results for kernel ridge regression. I obtain $L_{2}$ rates using integral operator techniques for kernel methods, drawing on [Smale and Zhou, 2007, Caponnetto and De Vito, 2007, Fischer and Steinwart, 2020], among others. See [Fischer and Steinwart, 2020] for references and a review of integral operator techniques and empirical process techniques used to analyze RKHS estimators in various norms, many of which rely on the effective dimension. My characterization of the counterfactual effective dimension appears to be new, and it avoids auxiliary approximation assumptions. The closed form solution of KRRR echoes classical solutions [Kimeldorf and Wahba, 1971, Schölkopf et al., 2001].
@@
I also build on the insight from semiparametric theory that the balancing weights for the average treatment effect are a special case of Riesz representers for causal parameters, which can be directly estimated [Robins et al., 2007, Avagyan and Vansteelandt, 2021, Chernozhukov et al., 2022b, Hirshberg and Wager, 2021, Chernozhukov et al., 2022a, Smucler et al., 2019, Hirshberg et al., 2019, Chernozhukov et al., 2020]. Similar to [Chernozhukov et al., 2022b, Chernozhukov et al., 2022a, Chernozhukov et al., 2020], I provide population $L_{2}$ rates. A key point of departure is that I study KRRR, which is a different estimator more closely aligned with kernel ridge regression, classical kernel ridge balancing weights, and popular R packages. Another distinction is that I place standard RKHS assumptions, which facilitates comparison with known optimality theory, including rate-optimal regularization. Finally, I avoid some auxiliary approximation assumptions of those works by defining and analyzing the counterfactual effective dimension; see Remark 5.
@@
The contribution of this work is not to provide new semiparametric theory. Rather, I demonstrate that KRRR has good properties similar to kernel ridge regression that were not fully articulated, and hence it is compatible with strong existing results, e.g. [Zheng and van der Laan, 2011, Chernozhukov et al., 2018, Rotnitzky et al., 2021, Kennedy, 2023], among many others. These strong results allow for some mis-specification and rate-optimal regularization, rather than the correct specification and undersmoothing studied by e.g. [Hirshberg et al., 2019, Mou et al., 2023]. See e.g. [Hirshberg and Wager, 2021, Ben-Michael et al., 2021] for detailed citations on balancing weights outside of the RKHS. See e.g. [van der Laan and Rose, 2011] for detailed citations on semiparametrics.
@@
An initial draft was circulated in February 2021 [Singh, 2021]. Prior to the current draft, [Wang et al., 2021] extend the kernel balancing weights of [Wong and Chan, 2018] to heterogeneous treatment effects and prove consistency under correct specification, without formal inference, a closed form, or $L_{2}$ rates for balancing weights. This paper’s results are complementary: formal inference, some mis-specification, a closed form, and $L_{2}$ rates for balancing weights, within a framework that includes other estimands. [^3] By formal inference, I mean justification for Gaussian approximation and nominal coverage guarantees. Subsequent works extend kernel balancing weights to confounding bridges [Kallus et al., 2021, Ghassami et al., 2021] and study other nonlinear function spaces [Chernozhukov et al., 2021].
@@ 3.1   Warm up: Balancing weights
These two formulations suggest estimation approaches. One option is to estimate the regression as $\hat{\gamma}$ then the ATE as $\mathbb{E}_{n}\{\hat{\gamma}(1,X)-\hat{\gamma}(0,X)\}$, where $\mathbb{E}_{n}(\cdot)$ is the empirical mean. Another is to estimate the balancing weight as $\hat{\alpha}$ then the ATE as $\mathbb{E}_{n}\{Y\hat{\alpha}(D,X)\}$. One may also combine these estimators into a so-called doubly robust procedure $\mathbb{E}_{n}[\hat{\gamma}(1,X)-\hat{\gamma}(0,X)+\hat{\alpha}(D,X)\{Y-\hat{\gamma}(D,X)\}]$. Each approach also has a sample splitting variant, where $(\hat{\gamma},\hat{\alpha})$ are estimated from, say $n/2$ held out observations, and the empirical mean is taken with respect to the remaining $n/2$ observations. For the sample splitting variant, $\hat{\alpha}$ must be a function that can be evaluated out-of-sample.
@@
Previous work proposes a kernel ridge balancing weight $\tilde{\alpha}$. My main contribution is to analyze its generalization error $\|\tilde{\alpha}-\alpha_{0}\|_{2}$, achieving the well-known lower bound for regression in the RKHS framework, without auxiliary approximation assumptions. Once a population $L_{2}$ rate is articulated, general semiparametric guarantees from the literature become available. These guarantees justify Gaussian approximation for causal functions and some mis-specification, unlike many previous works on kernel ridge balancing weights.
@@ 3.2   Generalizing from balancing weights to Riesz representers
The ATE is a special case of a causal parameter. This paper considers parameters of the form $\theta_{0}=\mathbb{E}\{m(W,\gamma_{0})\}$, where $W$ concatenates observed variables, $\gamma_{0}(W)=\mathbb{E}(Y|W)$ is a regression, and the formula $m$ accommodates various parameters beyond the ATE. For ATE, $W=(D,X)$ and $m(w,f)=f(1,x)-f(0,x)$. I study a class defined by a regularity condition generalizing the idea that the propensity score is away from zero and one.
Assumption 1 (Mean square continuity).
@@
Nonasymptotic analysis allows $\bar{M}$ to be a diverging sequence as the sample size increases. In particular, this class includes functionals of the form $f\mapsto\mathbb{E}\{m_{h}(W,f)\}$ where $m_{h}(w,f)=\ell_{h}(w)\check{m}(w,f)$. The weighting $\ell_{h}(w)$ is indexed by a vanishing bandwidth $h\downarrow 0$, which also indexes $\bar{M}=\bar{M}_{h}\uparrow\infty$. In this thought experiment, a sequence of semiparametric quantities, each satisfying Assumption 1, provides a pointwise approximation to a nonparametric causal function evaluated at a point. I provide two concrete examples.
Example 1 (Heterogeneous policy effects).
Here $W=X$ are regressors and $\check{m}(w,f)=f\{t(x)\}-f(x)$, where $t$ is a known mapping that serves as the counterfactual policy of interest. The weighting $\ell_{h}(w)=\textsc{local}\{(x_{1}-x_{1}^{*})/h\}$ is a localization around the location $X_{1}=x_{1}^{*}$ (see Appendix A for details). Suppose that the density ratio is bounded above, i.e. $\omega(W)=\textsc{density}\{t(W)\}/\textsc{density}(W)\leq\bar{\omega}$ almost surely. Suppose that the weighting function at a fixed $h$ is bounded, i.e. $|\ell_{h}(W)|\leq\bar{\ell}_{h}$ almost surely. Then $\bar{M}_{h}=2\bar{\ell}_{h}^{2}(\bar{\omega}+1)$.
Example 2 (Heterogeneous treatment effects).
@@
For further interpretation, suppose the the first covariate $X_{1}$ is continuous. In the finite sample, Example 2 is the heterogeneous treatment effect for the subpopulation whose first covariate is local to $x_{1}^{*}$. As $h\downarrow 0$, Example 2 converges to the heterogeneous treatment effect for the subpopulation whose first covariate value is $x_{1}^{*}$. Section 7 gives a real world example where formal uncertainty quantification leads to economic insights.
Taking $\ell_{h}(w)=1$, Examples 1 and 2 reduce to the average policy effect and average treatment effect, respectively. The class defined by Assumption 1 contains not only average effects but also heterogeneous effects, i.e. pointwise approximations of causal functions.
The balancing weight is a special case of a Riesz representer. It is well known that each parameter within the class defined by Assumption 1 has both a regression formulation and a generalized balancing weight formulation.
Lemma 1 (Riesz representation; Lemma S3.1 of [Chernozhukov et al., 2022b]).
Suppose Assumption 1 holds and $\gamma_{0}\in\mathcal{G}\subset L_{2}$. Then there exists a Riesz representer $\alpha_{0}\in L_{2}$ such that $\mathbb{E}\{m(W,f)\}=\mathbb{E}\{\alpha_{0}(W)f(W)\}$ for all $f\in\mathcal{G}$. Moreover, there exists a unique minimal Riesz representer $\alpha_{0}^{\min}\in closure\{span(\mathcal{G})\}$ that satisfies this equation.
For the rest of the paper, I will set aside the “correct specification” assumption that $\gamma_{0}\in\mathcal{G}$, where $\mathcal{G}$ is a known subset of $L_{2}$ that can be imposed in estimation of $\hat{\gamma}$. As such, there exists a unique Riesz representer $\alpha_{0}=\alpha_{0}^{\min}$. See Remark 3 for a list of many works that place “correct specification” style assumptions requiring $\gamma_{0}$ to be within the RKHS.
For a functional $f\mapsto\mathbb{E}\{m_{h}(W,f)\}$ where $m_{h}(w,f)=\ell_{h}(w)\check{m}(w,f)$, its Riesz representer is $\alpha_{h}(w)=\ell_{h}(w)\check{\alpha}_{0}(w)$, where $\check{\alpha}_{0}(w)$ is the Riesz representer for $f\mapsto\mathbb{E}\{\check{m}(W,f)\}$.
In summary, previous works on classical kernel ridge balancing weights provide formal inference for a limited class of causal parameters, i.e. average effects. They also typically require correct specification of $\gamma_{0}$ within the RKHS. I study the broader class defined in Assumption 1, which includes heterogeneous policy effects (Example 1) and heterogeneous treatment effects (Example 2). I take the agnostic stance of $\gamma_{0}\in L_{2}$. By analyzing generalization error $\|\tilde{\alpha}-\alpha_{0}\|_{2}$, general semiparametric guarantees become immediate.